OpenStudy (anonymous):

The diameter of the larger circle is 12.5 cm. The diameter of the smaller circle is 3.5 cm. What is the approximate area of the shaded region? A. 28.26 cm² B. 113.04 cm² C. 132.27 cm² D. 254.34 cm²

OpenStudy (anonymous):

@amberosales @fbi2015 @freegirl112 @iGreen

OpenStudy (igreen):

Find the areas of the two circles and subtract.

OpenStudy (igreen):

\(A = \pi r^2\)

OpenStudy (igreen):

Divide the diameters by 2 to find the radius's. 12.5 / 2 = 6.25 3.5 / 2 = 1.75

OpenStudy (igreen):

Larger circle: \(A = (3.14)(6.25^2)\) Smaller circle: \(A = (3.14)(1.75^2)\) Simplify and subtract.

OpenStudy (anonymous):

so it would be b? @iGreen

OpenStudy (anonymous):

@iGreen @iGreen @iGreen

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

thanks :) @iGreen

OpenStudy (igreen):

No problem, happy to help \(\Huge\ddot\smile\)